OpenStudy (anonymous):

Find the quotient: x^2-16/x+4...... would this be the answer? x^2-16/x+4= 2+4= 6

OpenStudy (lgbasallote):

how did you get 2 + 4 = 6? i know 2 + 4 is 6...but what is its relation to the problem?

OpenStudy (anonymous):

x^2-16/x+4= 6 ?

OpenStudy (anonymous):

go to mathway.com and type it in >.>

jimthompson5910 (jim_thompson5910):

Hint: x^2 - 16 = x^2 - 4^2 x^2 - 16 = (x - 4)(x + 4)

OpenStudy (anonymous):

(x - 4)(x + 4)/(x+4)